# On the Fontaine-Mazur conjecture for $p=3$

### v3

@@ Abstract.
In this article, we prove some new cases of the Fontaine-Mazur conjecture on two-dimensional Galois representations when $p=3$. Combining the previous work, we can conclude the conjecture in the regular case when $p$ is an odd prime. Our result is a sequel to Pan’s work, based on some recent progress on $p$-adic Langlands correspondence and Galois deformation theory.
@@ 1. Introduction
We will follow the strategy of Pan’s proof. However, Pan’s assumption is frequently used in his paper [Pan22]. Roughly speaking, the utilization of it is mainly in the following aspects:
  • local-global compatibilty,
  • Galois deformation theory,
  • finding “nice” primes.
@@
For the first part, it is an application of Paškūnas theory. In [Pan22], Pan uses the results in [Paš13], which excludes the case $p=3$ somewhere. Recently, the missing case has been studied in [PT21] (even for $p=2$). Based on the results for small primes, Tung studied the Fontaine-Mazur conjecture under Talylor-Wiles hypothesis for $p=2$ in [Tun21b] and $p=3$ in [Tun21a]. In this article, we may also use these results to obtain the same local-global compatibility arguments (see Section 3.3) and further prove some $R=\mathbb{T}$ theorems (see Theorem 4.1.4 and Theorem 4.3.3) without Pan’s assumption with almost the same proof as Pan’s, replacing the old work by the new one.
For the second part, the universal (pseudo-)deformation ring seems to be more difficult to study in the residually reducible case without Pan’s assumption. For a finite character $\bar{\chi}:G_{\mathbb{Q}_{p}}\to\mathbb{F}^{\times}$, if $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=1$ (for example, $\bar{\chi}\neq\mathbf{1},\omega^{-1}$), then by [Kis09, Corollary 1.4.4], there is a natural surjection from the universal pseudo-deformation ring to the universal (unframed) deformation ring. Here $\mathbb{F}$ is a finite field of characteristic $p$. (If $p\geq 5$, then after a finite twist, we may assume $\bar{\chi}\neq\omega^{-1}$.) Using this surjection, the local pseudo-deformation ring (at $p$) is well-understood (see [Paš13, Appendix B] for details). In [Pan22], an explicit description of the local pseudo-deformation ring seems to be essential so that we can study the ordinary locus of the global universal pseudo-deformation ring. See [Pan22, Lemma 7.4.4 & 7.4.8].
In the case $p=3$ and $\bar{\chi}=\omega^{-1}$ (hence quadratic), by the local Euler characteristic formula, we have $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=2$. Then the arguments above do not hold, and the local pseudo-deformation ring is unknown (at least for the author). In spite of this, some geometric properties of the local pseudo-deformation ring have been studied in the recent work [BIP23] and [BJ23], in which settings are more general. These are enough for us to establish analogous results to Pan’s.
Another obstruction is the global deformation theory. Consider a finite character $\bar{\chi}:G_{F}\to\mathbb{F}^{\times}$, and a non-split extension $\bar{\rho}_{b}$ of $\mathbf{1}$ by $\bar{\chi}$, where $F$ is an abelian totally real field. If $H^{0}(G_{F},\operatorname{ad}^{0}(\bar{\rho}_{b}(1)))=0$ (for example $\bar{\chi}\neq\omega^{-1}$), then the universal unframed ordinary deformation ring $R_{b}^{\mathrm{ord}}$ is of dimension at least $[F:\mathbb{Q}]+1$, which is the same as dimension of the Iwasawa algebra $\Lambda_{F}$. (Again, if $p\geq 5$, then after a finite twist, we may assume $\bar{\chi}\neq\omega^{-1}$.) Using Pan’s finiteness result in the ordinary case ([Pan22, Theorem 5.1.1]), we can find enough pro-modular primes in the ordinary locus of the global pseudo-deformation ring.
If $p=3$ and $\bar{\chi}=\omega^{-1}$, then the lower bound of the dimension of $R_{b}^{\mathrm{ord}}$ is $[F:\mathbb{Q}]$ (see [Pan22, Lemma 5.4.2]). In this situation, we may not find pro-modular primes in the ordinary locus. To overcome this difficulty, we need a generalized Greenberg-Wiles formula (Proposition 2.4.3) to study the dimension of some irreducible component of $R_{b}^{\mathrm{ord}}$, which is one of the main innovations of this paper.
@@
For the third part, combining the $R=\mathbb{T}$ result (Theorem 4.1.4), to prove our main theorem in the residually reducible case, we just need to find a “potentially nice” prime (see [Pan22, 7.2.1]) in each irreducible component of the spectrum of the universal pseudo-deformation ring $R^{\mathrm{ps}}$. However, in the case $p=3$ and $\bar{\chi}=\omega$ (hence quadratic), it seems to be difficult to find a nice prime in some situation as we cannot ignore the “dihedral condition” (see 3) of Definition 4.1.3 (2) and Lemma 4.1.7). The strategy to solve this problem is to use the method developed in [Zha24] so that we only need to find an irreducible one-dimensional pro-modular prime in each irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}$ of large dimension, which is much more convenient to find out. Furthermore, it allows us to prove a big $R=\mathbb{T}$ theorem in the residually reducible case in fact.
@@ 1.1. Notations
For a prime ideal $\mathfrak{p}$ of a commutative ring $R$, we denote its residue field by $k(\mathfrak{p})$. Let $R_{\mathfrak{p}}$ be the localization at $\mathfrak{p}$. We write $\hat{R_{\mathfrak{p}}}$ as its $\mathfrak{p}$-adic completion. We say a commutative ring $R$ is a CNL ring if it is a complete noetherian local ring, and we usually denote its maximal ideal by $\mathfrak{m}_{R}$.
@@ 2.2. Pseudo-representations
2) there exists an order 2 element $c\in G$ such that $T(c)=0$,
3) $T(\sigma\tau)=T(\tau\sigma)$ for all $\sigma,\tau\in G$,
4) $T(\gamma\delta\eta)+T(\gamma\eta\delta)-T(\gamma\eta)T(\delta)-T(\eta\delta)T(\gamma)-T(\delta\gamma)T(\eta)+T(\gamma)T(\delta)T(\eta)=0,$ for any $\delta,\gamma,\eta\in G$.
The determinant $\operatorname{det}(T)$ of a pseudo-representation $T$ is a character (using 4)) defined by
@@
For a pseudo-representation $T$ and $\sigma,\tau\in G$, define:
@@ 2.3. Local pseudo-deformation rings
In this subsection, we collect some results about the universal pseudo-deformation rings.
@@ 2.4. Galois cohomology
Let $p$ be a prime. Let $K$ be a local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$. Let $G$ be a profinite group, and $V$ a finite dimensional $K$-vector space with continuous $G$-action. At first, we give the Euler-Poincaré characteristic formulae in the version for $V$. In the case that $V$ is a finite $G$-module, one can see [NSW08, Theorem 7.3.1] and [NSW08, 8.7.4].
@@
For the case $l=p$, one can see [BJ23, Theorem 3.4.1]. For the case $l\neq p$, the proof is similar. ∎
@@ 2.5. Global Galois deformation theory
Let $R^{\mathrm{ps}}$ be the universal pseudo-deformation ring parametrizing all liftings of the pseudo-representation $1+\bar{\chi}$ with determinant $\chi$. Let $\mathfrak{p}\in\operatorname{Spec}R^{\mathrm{ps}}$ of characteristic 00 and suppose that it defines an irreducible pseudo-representation. Then by Definition 2.2.3, we can construct an absolutely irreducible Galois representation $\rho(\mathfrak{p})^{\circ}:G_{F,T}\to\operatorname{GL}_{2}(A)$, where $A$ is the normal closure of $R^{\mathrm{ps}}/\mathfrak{p}$. Write $\bar{\rho}_{b}=\rho(\mathfrak{p})^{\circ}\mod\mathfrak{m}_{A}$, where $0\neq b\in H^{1}(G_{F,T},\mathbb{F}(\bar{\chi}^{-1}))$. Then by [SW99, Lemma 2.13], up to some non-zero scalars of $\mathbb{F}$, $b$ is independent of the choice of $\rho(\mathfrak{p})^{\circ}$ given in Definition 2.2.3.
@@
In the rest of this subsection, we study the universal deformation ring of the Galois representation $\rho(\mathfrak{p}):=\rho(\mathfrak{p})^{\circ}\otimes k(\mathfrak{p})$. Suppose that $\rho(\mathfrak{p})$ is ordinary at each $v|p$.
@@ 3.1. Completed cohomology
$S_{\psi}(U^{p}):=\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(K/\mathcal{O},S_{\psi}(U^{p},K/\mathcal{O}))$
@@
$M_{\psi}(U^{p}):=S_{\psi}(U^{p},K/\mathcal{O})^{\vee}=\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(S_{\psi}(U^{p},K/\mathcal{O}),K/\mathcal{O})$
@@
Hence $S_{\psi}(U^{p},K/\mathcal{O})$ is a smooth admissible $\mathcal{O}$-representation of $D_{p}^{\times}$. It is also clear that $S_{\psi}(U^{p})$ is $p$-torsion free and
@@ 3.3. Local-global compatibility
    The semi-simple Weil-Deligne representation $\mathrm{WD}(\varepsilon\otimes\rho(\mathfrak{p})_{v})^{\mathrm{ss}}$ corresponds to $\pi_{v}^{K(\mathfrak{p})}$ under the Hecke correspondence. In our case, $\mathrm{WD}(\varepsilon\otimes\rho(\mathfrak{p})_{v})$ is just $(\varepsilon\otimes\rho(\mathfrak{p})_{v}\otimes B_{\mathrm{cris}})^{G_{F_{v}}}$ with $\operatorname{Frob}_{v}$ acting via $\varphi$.
@@
Let $\Pi_{v}$ be the universal unitary completion of $(\operatorname{Sym}^{k_{\sigma_{v}-2}}(K^{2})\otimes\det{}^{w_{\sigma_{v}}})^{*}\otimes\pi^{K(\mathfrak{p})}_{v}$ as a $K$-representation of $\mathrm{GL}_{2}(F_{v})$. Note that $\pi^{K(\mathfrak{p})}_{v}$ is an irreducible principal series as the level is $\mathrm{GL}_{2}(O_{F_{v}})$ (hence unramified). Otherwise it is one-dimensional and $\pi_{\mathfrak{p}}$ has to factor through the reduced norm map by the approximation Theorem (for example, see [Gee22, 4.17] or [KST20, Lemma 6.2] in general). Then $\Pi_{v}$ is a topologically irreducible admissible unitary $\mathrm{GL}_{2}(F_{v})$-representation. Here we use [Paš09, Theorem 1.1] and [BB10, Théorème] for the non-ordinary case and [BE10, Proposition 2.2.1] for the ordinary case.
@@
Let $\Pi_{v}^{0}$ be a $\mathrm{GL}_{2}(F_{v})$-invariant bounded open ball of $\Pi_{v}$ and let $(\Pi_{v}^{0})^{d}=\operatorname{Hom}_{\mathcal{O}}^{\mathrm{cont}}(\Pi_{v}^{0},\mathcal{O})$ be its Schikhof dual. Then we also have

### v4

@@ Abstract.
In this article, we prove the remaining open cases of the Fontaine-Mazur conjecture on two-dimensional regular Galois representations over $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ when $p=3$, hence concluding the conjecture in the regular case for all odd primes. Our result is a sequel to Pan’s work, based on some recent progress on $p$-adic Langlands correspondence, Galois deformation theory and a potential pro-modularity result.
@@ 1. Introduction
The main difficulty of the Fontaine-Mazur conjecture at $p=3$ is the residually reducible case. To outline the proof of Theorem 1.0.2 in this case, we suppose the semisimplification of $\bar{\rho}$ is (up to twist) isomorphic to $1\oplus\bar{\chi}$ for some continuous odd almost unramified character $\bar{\chi}:G_{\mathbb{Q}}\to\mathbb{F}^{\times}$, where $\mathbb{F}$ is a finite field of characteristic $p$.
Our strategy is a refinement of Pan’s. Precisely, we prove a potential pro-modularity argument, rather than just identify some irreducible component of the universal deformation ring as the in the big Hecke algebra, so that our strategy works for all odd primes.
To illustrate the obstructions in the case $p=3$ and $\bar{\chi}|_{G_{\mathbb{Q}_{p}}}=\omega^{-1}$ (we call it “exceptional” afterwards) in Pan’s method, we first briefly recall the main steps of Theorem 1.0.1 in the residually reducible case.
1) Local-global compatibilty results. There are two main purposes in this part. The first one is to justify whether the Galois representation defined by a map from the big Hecke algebra to $\overline{\mathbb{Q}_{p}}$ comes from a cuspidal automorphic representation. The second one is to collect some necessary results for the patching argument (see [Pan22, page 1033] for details). Most results in this part are applications of Paškūnas theory [Paš13], which excludes our exceptional case.
2) Patching at nice primes. In this part, Pan proves an $R_{\mathfrak{q}}=\mathbb{T}_{\mathfrak{q}}$ theorem following the strategy of Skinner-Wiles [SW99], where $\mathfrak{q}$ is a one-dimensional prime (called “nice”) of the big Hecke algebra satisfying some special properties. In other words, we can identify some irreducible components of the universal pseudo-deformation ring as in the big Hecke algebra, if those contain a “nice” prime.
3) Ordinary Fontaine-Mazur conjecture in the residually reducible case. In this part, Pan generalizes the main theorem of [SW99] and proves a finiteness result between the Iwasawa algebra and the universal ordinary pseudo-deformation ring. This allows us to find enough pro-modular primes, i.e. primes coming from the big Hecke algebra, in the ordinary locus of the universal pseudo-deformation ring.
4) Galois deformation theory for pseudo-deformation rings. From the previous three steps, we only need to show that there exists an irreducible component of $\operatorname{Spec}R$ containing both a “nice” prime and the one defined by our Galois representation $\rho$. To produce enough “nice” primes in the ordinary locus of $\operatorname{Spec}R$ (write as $\operatorname{Spec}R^{\operatorname{ord}}$), we first need to know the kernel from the local pseudo-deformation at $p$ to its ordinary locus (write the map as $R_{p}\to R_{p}^{\operatorname{ord}}$), and obtain an exact bound of the Krull dimension of $\operatorname{Spec}R^{\operatorname{ord}}$. In this step, Pan only finds “potentially nice” primes, which is enough for his purpose.
Here we point out some differences and difficulties in the exceptional case and describe the improvements in this paper:
@@
In step 1), as we mentioned before, some results in the exceptional case are missing in [Paš13]. The reason is that the behavior of the block in this case is specialzed and more complicated (see the list in A.2). Recently, this case has been studied in [PT21] (even for $p=2$). Based on Paškūnas theory for small primes, Tung studied the Fontaine-Mazur conjecture for $p=2$ in [Tun21b] and $p=3$ in [Tun21a] under Talylor-Wiles hypothesis. In this article, we also use these results to obtain the same local-global compatibility arguments (see Section 3.3) and further prove some $R=\mathbb{T}$ theorems (see Theorem 4.1.4 and Theorem 4.3.3) without Pan’s assumption.
In step 2), some crucial properties of the local framed deformation ring at $p$, such as its normality, for proving the patching argument were missing at the exceptional case. Combining recent progress [BIP23] on local framed deformation ring at $p$, we are able to fill in such gaps in Pan’s strategy.
In step 4), there are several obstructions, and here for instance, we explain the most subtle case, $p=3$ and $\bar{\chi}=\omega=\omega^{-1}$.
  • When $p\geq 5$ and $\bar{\chi}|_{G_{\mathbb{Q}_{p}}}\neq\mathbf{1},\omega^{-1}$ (after a twist, we may assume this as $\omega$ is not quadratic), the kernel of $R_{p}\to R_{p}^{\operatorname{ord}}$ has been studied in [Paš13, Appendix B]. Note that in this case, we have $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=1$. Then by [Kis09, Corollary 1.4.4], there is a natural surjection (hence isomorphism) from the universal pseudo-deformation ring to the universal (unframed) deformation ring. Thus, we can calculate the pseudo-deformation ring explicitly. See [Pan22, Lemma 7.4.4 & 7.4.8]. However, such a calculation is unknown when $p=3$ and $\bar{\chi}=\omega^{-1}$ as $\dim_{\mathbb{F}}H^{1}(G_{\mathbb{Q}_{p}},\mathbb{F}(\bar{\chi}^{-1}))=2$. Despite the difficulties, combining some geometric properties of the local pseudo-deformation ring studied in the recent work [BIP23] and [BJ23], we are able to develop analogous results to tackle this issue without knowing the structure of the local pseudo-deformation ring.
  • When $p\geq 5$ and $\bar{\chi}\neq\omega^{-1}$ (again, we may assume this as $\omega$ is not quadratic), then we have $H^{0}(G_{\mathbb{Q}},\operatorname{ad}^{0}(\bar{\rho}_{b}(1)))=0$. Then by the Greenberg-Wiles formula, the universal unframed ordinary deformation ring $R_{b}^{\mathrm{ord}}$ is of dimension at least $2$, and hence of the same dimension as the Iwasawa algebra. By Pan’s finiteness result (in step 3)), we can find enough pro-modular primes in $\operatorname{Spec}R^{\operatorname{ord}}$, and also “potentially nice” primes. In our case, the dimension of $H^{0}$ is $1$, leading the lower bound of $R_{b}^{\mathrm{ord}}$ to be $1$. Hence, the arguments above cannot work. To overcome this difficulty, we prove a generalized Greenberg-Wiles formula (Proposition 2.4.3) to study the dimension of some irreducible component of $R_{b}^{\mathrm{ord}}$.
  • In the definition of a “nice” prime (see Definition 4.1.3), we need to know the corresponding representation is not dihedral, i.e. induced from a quadratic character. When $p\geq 5$ and $\bar{\chi}=\omega$, this condition is trivially satisfied, whereas in our case, we must not ignore it. Note that by Leopoldt’s conjecture for the abelian CM field $F(\mu_{3})$, the upper bound of the dihedral locus of $\operatorname{Spec}R$ is $[F:\mathbb{Q}]+2$, where $F$ is an appropriate abelian base change of $\mathbb{Q}$. Thus, to find (potentially) “nice” primes, we first need to find a pro-modular prime with larger dimension to avoid the dihedral locus. However, in the non-generic reducible case (see Section 5.4), the lower bound of the dimenson of the pro-modular prime we can find there is $[F:\mathbb{Q}]+2$, which is not large enough to find a (potentially) “nice” prime when we are in the exceptional case. Therefore, we need a new strategy of this part here. See Remark 5.4.8.
@@
    In this paper, we use the method developed in the author’s previous work [Zha24] so that we only need to find an irreducible one-dimensional pro-modular prime instead of a “nice” prime in each irreducible component of $\operatorname{Spec}R^{\mathrm{ps}}$ of large dimension, which is much more convenient to find out. Furthermore, it allows us to prove a potential big $R=\mathbb{T}$ theorem in the residually reducible case in fact. See Remark 5.1.3.
As we mentioned in the previous discussions, we actually prove the following potential pro-modularity result in the residually reducible case, which is the main innovation in this paper. To state it more clearly, we give the following settings.
Let $p$ be an odd prime and let $F$ be an abelian totally real field in which $p$ splits completely. Write $\Sigma_{p}$ as the set of places of $F$ above $p$ and let $\Sigma$ be a finite set of finite places of $F$ containing $\Sigma_{p}$. Let $\bar{\chi}:G_{F,\Sigma}\to\mathbb{F}^{\times}$ be a continuous, odd, almost unramified character which can be extended to a character of $G_{\mathbb{Q}}$. Then $T=1+\bar{\chi}$ defines a pseudo-representation. Further assume that for any $v|p$, $\bar{\chi}|_{G_{F_{v}}}\neq\mathbf{1}$. Let $\chi$ be a de Rham lifting of $\bar{\chi}$.
Theorem 1.0.3 (Theorem 5.5.1, Proposition 5.5.2).
There exists an abelian totally real extension $F_{1}/F$ of even degree such that $p$ splits completely in $F_{1}$ and for any irreducible component of the universal pseudo-deformation ring $R_{F_{1}}^{\mathrm{ps}}$ (parametrizing all pseudo-deformations of $T|_{G_{F_{1}}}$ unramified out places lying above $\Sigma$ with fixed determinant $\chi$) of dimension at least $1+2[F_{1}:\mathbb{Q}]$, it is pro-modular of dimension $1+2[F_{1}:\mathbb{Q}]$. The corresponding big Hecke algebra over $F_{1}$ is equidimensional of dimension $1+2[F_{1}:\mathbb{Q}]$.
Remark 1.0.4.
1) Actually, by Proposition 2.5.2, we know that for any irreducible component of $R_{F_{1}}^{\mathrm{ps}}$, it is either of dimension at least $1+2[F_{1}:\mathbb{Q}]$ or contained in its reducible locus, hence of dimension at most $2$ by Leopoldt’s conjecture. Thus, we may view this result as a potential big $R=\mathbb{T}$ theorem, identifying $\mathbb{T}$ as the part of $R$ with large dimension.
2) In a future work, we will discuss more about applications of this potential pro-modularity result.
@@ 1.1. Notations
For a prime ideal $\mathfrak{p}$ of a commutative ring $R$, we denote its residue field by $k(\mathfrak{p})$. Let $R_{\mathfrak{p}}$ be the localization at $\mathfrak{p}$. We write $\widehat{R_{\mathfrak{p}}}$ as its $\mathfrak{p}$-adic completion. We say a commutative ring $R$ is a CNL ring if it is a complete noetherian local ring, and we usually denote its maximal ideal by $\mathfrak{m}_{R}$.
@@ 2.2. Pseudo-representations
2) $T(\sigma\tau)=T(\tau\sigma)$ for all $\sigma,\tau\in G$,
3) $T(\gamma\delta\eta)+T(\gamma\eta\delta)-T(\gamma\eta)T(\delta)-T(\eta\delta)T(\gamma)-T(\delta\gamma)T(\eta)+T(\gamma)T(\delta)T(\eta)=0,$ for any $\delta,\gamma,\eta\in G$.
The determinant $\operatorname{det}(T)$ of a pseudo-representation $T$ is a character (using 3)) defined by
@@
If there exists an order $2$ element $c\in G$ such that $T(c)=0$, then for a pseudo-representation $T$ and $\sigma,\tau\in G$, we define:
@@ 2.3. Local pseudo-deformation rings
In this subsection, we collect some results about local universal pseudo-deformation rings.
@@ 2.4. Galois cohomology
Let $p$ be a prime. Let $K$ be a $p$-adic local field with its ring of integers $\mathcal{O}$, a uniformizer $\varpi$ and its residue field $\mathbb{F}$. Let $G$ be a profinite group, and $V$ a finite dimensional $K$-vector space with continuous $G$-action. At first, we give the Euler-Poincaré characteristic formulae in the version for $V$. In the case that $V$ is a finite $G$-module, one can see [NSW08, Theorem 7.3.1] and [NSW08, 8.7.4].
@@
See [BBD+13, page 78, Theorem 3.9.1]. For the case $l=p$ and $K$ a local field of characteristic $p$, this result is also true and one can see [BJ23, Theorem 3.4.1]. ∎
@@ 2.5. Global Galois deformation theory
Let $R^{\mathrm{ps}}$ be the universal pseudo-deformation ring parametrizing all liftings of the pseudo-representation $1+\bar{\chi}$ with determinant $\chi$. Let $\mathfrak{p}\in\operatorname{Spec}R^{\mathrm{ps}}$ (we do not require $\varpi\notin\mathfrak{p}$ here) and suppose that it defines an irreducible pseudo-representation. Then by Definition 2.2.3, we can construct an absolutely irreducible Galois representation $\rho(\mathfrak{p})^{\circ}:G_{F,T}\to\operatorname{GL}_{2}(A)$, where $A$ is the normal closure of $R^{\mathrm{ps}}/\mathfrak{p}$. Write $\bar{\rho}_{b}=\rho(\mathfrak{p})^{\circ}\mod\mathfrak{m}_{A}$, where $0\neq b\in H^{1}(G_{F,T},\mathbb{F}(\bar{\chi}^{-1}))$. Then by [SW99, Lemma 2.13], up to some non-zero scalars of $\mathbb{F}$, $b$ is independent of the choice of $\rho(\mathfrak{p})^{\circ}$ given in Definition 2.2.3.
@@
In the rest of this subsection, we study the universal deformation ring of the Galois representation $\rho(\mathfrak{p}):=\rho(\mathfrak{p})^{\circ}\otimes k(\mathfrak{p})$. Suppose that $\rho(\mathfrak{p})$ is ordinary at each $v|p$ and $\varpi\notin\mathfrak{p}$.
@@ 3.1. Completed cohomology
$S_{\psi}(U^{p}):=\operatorname{Hom}_{\mathcal{O}}(K/\mathcal{O},S_{\psi}(U^{p},K/\mathcal{O}))$
@@
$M_{\psi}(U^{p}):=S_{\psi}(U^{p},K/\mathcal{O})^{\vee}=\operatorname{Hom}_{\mathcal{O}}(S_{\psi}(U^{p},K/\mathcal{O}),K/\mathcal{O})$
@@
Hence $S_{\psi}(U^{p},K/\mathcal{O})$ is a smooth admissible $\mathcal{O}$-representation of $D_{p}^{\times}$. It is also clear that $S_{\psi}(U^{p})$ is $\varpi$-torsion free and
@@ 3.3. Local-global compatibility
    The semi-simple Weil-Deligne representation $\mathrm{WD}(\varepsilon\otimes\rho(\mathfrak{p})_{v})^{\mathrm{ss}}$ corresponds to $\pi_{v}^{K(\mathfrak{p})}$ under the Hecke correspondence.
@@
Let $\Pi_{v}$ be the universal unitary completion of $(\operatorname{Sym}^{k_{\sigma_{v}-2}}(K^{2})\otimes\det{}^{w_{\sigma_{v}}})^{*}\otimes\pi^{K(\mathfrak{p})}_{v}$ as a $K$-representation of $\mathrm{GL}_{2}(F_{v})$. Note that $\pi^{K(\mathfrak{p})}_{v}$ is an irreducible principal series as the level is $\mathrm{GL}_{2}(O_{F_{v}})$ (hence unramified). Otherwise it is one-dimensional and $\pi_{\mathfrak{p}}$ has to factor through the reduced norm map by the approximation Theorem (for example, see [Gee22, Fact 4.17] or [KST20, Lemma 6.2] in general). Then $\Pi_{v}$ is a topologically irreducible admissible unitary $\mathrm{GL}_{2}(F_{v})$-representation. Here we use [Paš09, Theorem 1.1] and [BB10, Théorème] for the non-ordinary case and [BE10, Proposition 2.2.1] for the ordinary case.
@@
Let $\Pi_{v}^{0}$ be a $\mathrm{GL}_{2}(F_{v})$-invariant bounded open ball of $\Pi_{v}$ and let $(\Pi_{v}^{0})^{d}=\operatorname{Hom}_{\mathcal{O}}(\Pi_{v}^{0},\mathcal{O})$ be its Schikhof dual. Then we also have